RC meeting and Run Plan, Saturday February 4th
From SBS wiki
Return to the RC Meeting Calendar
- RC: Rachel Montgomery, firstname.lastname@example.org
- On-going production running with polhe3 and 45uA.
- Since yesterday morning shift crews have collected of 2.5C of uncorrected charge (3.1C of corrected charge).
- Current target polarisation is 49.4%, and it hit 52.3% yesterday evening after ~24hrs of production running.
- Production running polhe3.
- Continue production running at 45uA and polHe3
Data Collection and Polarisaton Status
- Accumulated charge uncorrected 22.7C; accumulated charge corrected 24.58C
- Current target polarisation 49.4%
Other Topics/Status Updates
- MCC are talking about recovery day Tue next week and accelerator tuning Thursday next week. Not cemented plans, will update when hear more.
- Shift Schedule Page
- Shift Opening: TO Sunday Feb 5th swing, please consider signing up
Requested Accesses to the Hall
- Looking to put SBS GEMs in DAQ during day shifts, need to fix a cable in Hall first
- Target team for oven heater check
- DC power did not access Hall during maintenance day for moeller magnet, would be good to coordinate in the future.
Notices for Shift Crew
- Check the How To page
- Pay attention to beam restoration procedures in the How To
- there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning or something which may change beam profile. Communicate with MCC to figure out why beam has been down and to figure out which kind of restoration you should do.
- production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
- NMR every 4 hours
- Note the instructions for performing an NMR if there has been any changes to magnetic fields - see https://logbooks.jlab.org/entry/4124321
- Contact RC any time for any questions